OpenStudy (silverlark2):

show that sq root 3 minus sq root 2 is equal to 1/sq root 3 plus sq root 2

OpenStudy (anonymous):

\[\frac{( \sqrt{3}-\sqrt{2}) }{ 1 }\]

OpenStudy (anonymous):

take conjugate

OpenStudy (anonymous):

\[ \frac{ [\sqrt(3)-\sqrt(2)][\sqrt(3) +\sqrt(2)] } { \sqrt(3) +\sqrt(2) }\]

OpenStudy (anonymous):